OpenStudy (anonymous):

What is the approximate volume of a cone with a height of 9 meters and base radius of 6 meters?

OpenStudy (anonymous):

You use the formula for a cone which is \[V = (1/3) \times \pi \times r ^{2} \times h\] Then you just plug in the values: \[V = (1/3) \times \pi \times 6 ^{2} \times 9\] \[V = 339.29 meters ^{3}\] Of course, that is a rounded answer.